Diffusion Moprning Report 11/16/10

Ted Berg tberg at stanford.edu
Tue Nov 16 06:49:56 PST 2010


Good morning All,
     Looking good this morning. Tylan LTO is down for pull and clean. 
Tylan 2 had MFS alarms. Possible temp fluctuations on Tylan 1. Epi 
testing after dome replacement. Ebeam still has computer issues. No 
other issues at this time.

					
	Date: 	11/16/10 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*yellow* 	*Possible temp fluctuations* 	** 	** 	*3*
Tylan 2 	*yellow* 	*intermittent mfs alarms and n2 flow issues* 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*up* 	** 	*.25/30* 	** 	**
Tylan 8-SiGe 	*up* 	** 	*20.1/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*12.8/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*16.86/30* 	** 	**
Tylan 11-LTO 	*down* 	*pull & clean* 	*0/30* 	** 	*1*
Tylan 12-FGA 	*up* 	*
* 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	** 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*yellow* 	*testing* 	** 	** 	**
AMAT Epi 	*yellow* 	*testing* 	** 	** 	*2*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*computer issues* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*A on line 633 PSI* 	*changeover working 
well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20101116/f103f184/attachment.html>


More information about the diffusion mailing list